OpenStudy (brucebaner):

What is the missing value? A. −13 B. −5 C. 5 D. 13

OpenStudy (danjs):

recall...

OpenStudy (danjs):

\[\frac{ 8^a }{ 8^b } = 8^{a - b}\]

OpenStudy (brucebaner):

can you help me @DanJS and @cwrw238

OpenStudy (danjs):

something + 4 = 9

OpenStudy (danjs):

a = not known b = -4 a - b = a-(-4) = a+ 4 = 9 a + 4 = 9

OpenStudy (brucebaner):

so that letter is it is it C

OpenStudy (danjs):

yep

OpenStudy (brucebaner):

thx

OpenStudy (brucebaner):

is it C @Michele_Laino

OpenStudy (danjs):

\[\frac{ 8^a }{ 8^{-4} } = 8^{a -(-4)} = 8^{a+4} = 8^9\]

OpenStudy (danjs):

since it is the same base, you can equate the powers a + 4 = 9

OpenStudy (brucebaner):

okay thx

OpenStudy (danjs):

just remember when you move a base to a power to the opposite side of a fraction... you switch the sign on the power, + to - or - to +

OpenStudy (michele_laino):

@brucebaner please follow the good reasoning of @DanJS , solve the equation a+4=9 for a

OpenStudy (brucebaner):

thx so much @Michele_Laino and @DanJS

OpenStudy (danjs):

your welcome, anytime

OpenStudy (michele_laino):

thanks! @brucebaner